### 14 Apr 2026 v2

# AQ-Stacker: An Adaptive Quantum Matrix Multiplication Algorithm with Scaling via Parallel Hadamard Stacking

Abstract
Matrix multiplication (MatMul) is the computational backbone of modern machine learning, yet its classical complexity remains a bottleneck for large-scale data processing. We propose a hybrid quantum-classical algorithm for matrix multiplication based on an adaptive configuration of Hadamard tests. By leveraging Quantum Random Access Memory (QRAM) for state preparation, we demonstrate that the complexity of computing the inner product of two vectors can be reduced to $O(\log N)$. We introduce an ”Adaptive Stacking” framework that allows the algorithm to dynamically reconfigure its execution pattern—from sequential horizontal stacking to massive vertical parallelism—based on available qubit resources. This flexibility enables a tunable time-complexity range, theoretically reaching $O(N^{2})$ on fault-tolerant systems while maintaining compatibility with near-term hardware. Our core theoretical contribution is the formalization of the ”Entropy Dividend”: an information-theoretic bound proving that the effective measurement variance $\sigma^{2}_{eff}\leq\frac{1-e^{-H(\psi)}}{S}$ suppresses sampling noise as the system entropy $H(\psi)$ increases. This makes AQ-Stacker uniquely suited for the stochastic weight distributions of deep neural networks. We validate the numerical stability of our approach through a Quantum Machine Learning (QML) simulation, achieving 96% accuracy on the MNIST handwritten digit dataset. Our results suggest that entropic noise suppression and parallel Hadamard stacking provide a scalable path toward super-classical efficiency in next-generation quantum-enhanced AI.
1 Introduction
Matrix multiplication (MatMul) is a cornerstone of modern computational science, forming the foundational layer for deep learning, scientific simulations, and large-scale data analysis [7]. While classical algorithms have reached an impressive $O(N^{2.37})$ complexity [1], the quadratic floor remains a formidable barrier for big-data applications. Quantum computing offers a theoretical path to sub-quadratic scaling, yet many proposed quantum linear algebra routines suffer from high circuit depths or rigid hardware requirements that render them ”galactic”—theoretically sound but practically unreachable for near-term devices.
In this work, we introduce AQ-Stacker (Adaptive Quantum Stacking), a hybrid algorithm that utilizes the Hadamard test as a primitive to perform matrix-vector and matrix-matrix multiplication. Central to our approach is the QRAM Input Model, which we treat as a standardized quantum-addressable data structure. By assuming the existence of a QRAM that provides state preparation in $O(\log N)$ time, we decouple the data-loading overhead-which typically scales as $O(N)$ for arbitrary state synthesis [12]—from the computational logic [6]. This allows us to focus on a novel architectural contribution: Adaptive Stacking.
AQ-Stacker addresses the ”Resource-Complexity Trade-off” by providing a tunable execution framework. Unlike static algorithms, AQ-Stacker can reconfigure its layout based on the available qubit width of the target processor. We demonstrate that by ”stacking” Hadamard tests vertically, the time complexity of an $N\times N$ MatMul can be reduced from the sequential $O(N^{3})$ to a parallelized $O(N^{2})$ (accounting for classical I/O) or even $O(\log N)$ in terms of pure quantum depth.
We validate our framework through a Quantum Machine Learning (QML) task, achieving a 96% classification accuracy on the MNIST dataset. This results proves that our adaptive MatMul approach maintains numerical stability and expressive power, making it a viable candidate for next-generation quantum-enhanced AI.
Core Contributions
The primary contributions of this work are summarized as follows:
  • AQ-Stacker Algorithm: A resource-adaptive hybrid algorithm that utilizes the Hadamard test as a computational primitive to perform matrix operations. By assuming an $O(\log N)$ QRAM interface, we decouple data-loading overhead from computational logic.
  • Adaptive Stacking Architecture: A flexible execution framework that reconfigures its layout (Horizontal, Balanced, or Vertical) based on available qubit width. We prove that vertical stacking can reduce quantum depth to $O(\log N)$ for $N\times N$ matrix multiplication.
  • The Entropy Dividend: We formalize a novel noise-suppression bound, $\sigma^{2}_{eff}\leq\frac{1-e^{-H(\psi)}}{S}$, proving that the high-entropy stochastic weights typical of neural networks inherently suppress measurement shot noise.
  • Numerical Stability Benchmarks: Through simulations on IRIS and MNIST, we demonstrate $96.0\%$ classification accuracy—a $90\%$ improvement over traditional Variational Quantum Circuits (VQC), confirming the viability of Hadamard stacking for deep learning.
2 Methodology
The AQ-Stacker algorithm decomposes matrix multiplication into three distinct phases: classical pre-processing, quantum execution via a resource-adaptive parallel Hadamard framework, and classical post-processing.
2.1 Mathematical Mapping: Classical to Quantum
To perform matrix multiplication via quantum inner products, we must map classical vectors in $\mathbb{R}^{N}$ to quantum states. Given two real-valued vectors $\vec{x}$ and $\vec{w}$, we define the mapping to normalized quantum states $|\psi\rangle$ and $|\phi\rangle$ as:
$|\psi\rangle=\frac{\vec{x}}{\|\vec{x}\|},\quad|\phi\rangle=\frac{\vec{w}}{\|\vec{w}\|}$ (1)
In a Real Euclidean space with an orthonormal basis, the classical dot product $\vec{x}\cdot\vec{w}=\sum_{i=1}^{n}x_{i}w_{i}$ is equivalent to the quantum inner product $\langle\psi|\phi\rangle$ scaled by the product of their Euclidean norms:
$\vec{x}\cdot\vec{w}=\langle\psi|\phi\rangle\cdot\|\vec{x}\|\cdot\|\vec{w}\|$ (2)
This equivalence allows AQ-Stacker to leverage the efficiency of the Hadamard test for overlap estimation while preserving the magnitude information of the original matrix entries through classical norm-tracking.
2.2 The Two-State Hadamard Test Primitive
The core computational unit of AQ-Stacker is the Hadamard test, used to estimate the inner product $\langle\phi|\psi\rangle$. Given two normalized vectors $v_{1},v_{2}\in\mathbb{R}^{N}$, we define the unitaries $U$ and $V$ such that $U|0\rangle^{\otimes n}=|\psi\rangle$ and $V|0\rangle^{\otimes n}=|\phi\rangle$. To compute the real part of the overlap, we initialize an ancilla qubit in the state $|0\rangle$ and the system register in $|\psi\rangle$.
The circuit evolution is as follows:
  1. 1.
    Superposition: Apply a Hadamard gate to the ancilla: $\frac{1}{\sqrt{2}}(|0\rangle+|1\rangle)\otimes|\psi\rangle$.
  2. 2.
    Controlled Operation: Apply a controlled-$U^{\dagger}V$ operation where $U^{\dagger}V$ acts on the system register only if the ancilla is $|1\rangle$: $\frac{1}{\sqrt{2}}(|0\rangle\otimes|\psi\rangle+|1\rangle\otimes U^{\dagger}V|\psi\rangle)$.
  3. 3.
    Interference: Apply a second Hadamard to the ancilla. The final state before measurement is:
    $|\Psi_{\text{final}}\rangle=\frac{1}{2}|0\rangle\otimes(I+U^{\dagger}V)|\psi\rangle+\frac{1}{2}|1\rangle\otimes(I-U^{\dagger}V)|\psi\rangle$ (3)
The probability of measuring the ancilla in state $|0\rangle$ is given by $P(0)=\frac{1}{2}(1+\text{Re}\langle\phi|\psi\rangle)$. After measurement, the inner product is given by $\langle\psi|\phi\rangle=P(0)-P(1)$, finally apply Equation 2 to reconstruct the dot product. Note that, in the context of machine learning, for real-valued vectors like MNIST, this suffices to calculate the matrix product $C=AB$ as the collection of dot products of the row-column tuples $C_{ij}=(Row_{i}A)\cdot(Column_{j}B)$.
2.3 Hadamard Test vs. other Quantum Objects
Our work makes the case for the Hadamard Test as the most efficient object for matrix multiplication in the context of machine learning, however it is not the only object we tested. The next section summarizes our comparisson results.
  • SWAP Test: it does not preserve negative phases, bacause it measures the absolute square of the inner product ($|\langle\psi|\phi\rangle|^{2}$), resulting in the loss of the specific sign (direction). In Machine learning, negative weights are essential for a model’s ability to learn complex patterns and represent a wide range of functions. The Hadamard test, in the other hand, measures the real part of the inner product while preserving the phase. Note that, for machine learning (where weights are real-valued) this is sufficient. A second circuit (to measure the imaginary part) would be needed for complex-valued matrices.
  • Block Encoding: A technique to embed a non-unitary matrix A into the top-left block of a larger unitary. In hardware, to reconstruct the full $N\times N$ matrix product classically, we need to perform full quantum state tomography, which is exponentially slow and usually defeats the purpose of the quantum speedup (Tomography Cost)[9]. Even in simulation, the excesive number of ancilla and data qubits can consume the memory as shown in Table 1
    Table 1: Resource usage in CPU/GPU simulation for the FABLE[2] block encoder.
    $D=2^{n}$ Data Qubits (n) Total Ancilla 2(n+1) Total Qubits CPU GPU
    2 1 4 5 0.008649 0.255813
    4 2 (IRIS) 6 8 0.008258 0.262382
    8 3 8 11 0.666807 0.330950
    16 4 10 14 352.36 2.734008
    32 5 12 17 ERROR: Insufficient memory Required: 262144M ERROR: Required memory: 262144M, max: 85480M (Host) + 40442M (GPU)
    1024 10 (MNIST) 22 32 OOM OOM
  • Linear Combination of Unitaries (LCUs): A powerful method for block-encoding a matrix A by decomposing it into a sum of unitaries
    $A=\sum_{j}\alpha_{j}U_{j}$
    where $U_{j}$ are unitary matrices and $\alpha_{j}$ are scalar coefficients. In the Pauli basis, the number of terms can grow exponentially for dense, random matrices reaching up to $4^{n}$ for an $n$-qubit system[3].
In summary, the Hadamard test achieves logarithmic storage capacity (the number of qubits decreases as the size of the input increases), a low circuit depth (compared to other objects), minimal number of ancilla, and preserves the phase of the inner product. It is highly scalable, flexible, and arguably a swiss-army knife for quantum computing.
2.4 Algorithmic Stages
Stage 1: Adaptive Circuit Preparation and Data Loading
AQ-STACKER utilizes a dual-path data loading architecture to ensure both immediate viability and long-term scalability.
  • Theoretical QRAM Interface: By assuming a standard Quantum Random Access Memory (QRAM) model, the algorithm achieves state preparation in $O(\log N)$ depth. This provides a theoretical path to a total quantum depth of $O(\log N)$ for a full $N\times N$ matrix-vector multiplication.
  • Near-Term Memoization Mode: In the absence of physical QRAM, we propose a classical memoization strategy to mitigate the traditional $O(N^{3})$ preparation bottleneck. By pre-calculating the isometric decompositions for the $2N$ unique vectors in $A$ and $B$, we reduce the heavy mathematical overhead of state synthesis from $O(N^{3})$ to $O(N^{2})$.
This hybrid approach ensures that the classical pre-processing cost strictly matches the fundamental $O(N^{2})$ I/O floor of matrix operations, making AQ-STACKER the first Hadamard-based framework that remains classically competitive while awaiting QRAM maturity.
In this stage, the algorithm prepares the metadata required for the quantum execution. Crucially, the classical Euclidean norms $\|A_{i}\|$ and $\|B_{j}\|$ are calculated for each of the $N$ rows of $A$ and $N$ columns of $B$. Since each norm calculation is $O(N)$, the total classical overhead for norm pre-calculation is $O(N^{2})$.
These norms are stored as metadata alongside the QRAM Input Model. We assume that row and column vectors are addressable in $O(\log N)$ time. The algorithm then generates a buffer of $N^{2}$ Hadamard test objects, each mapped to the normalized states $|\psi_{i}\rangle$ and $|\phi_{j}\rangle$. By pre-calculating norms at the matrix level rather than the element level, we ensure that the classical preparation phase remains strictly $O(N^{2})$, preserving the quadratic scaling of the overall framework.
Remark 1 (State Preparation Overhead).
While AQ-Stacker assumes an $O(\log N)$ QRAM interface for state preparation [6], we acknowledge that on near-term hardware lacking a physical QRAM, state synthesis for arbitrary vectors may scale as $O(N)$ [12]. In such regimes, the ”Turbo” mode depth remains $O(N)$, though the parallel Hadamard stacking still provides a linear speedup over sequential inner-product estimation.
Remark 2 (The QRAM-Bottleneck Mitigation Strategy).
While the theoretical $O(\log N)$ scaling relies on a quantum-addressable memory, the practical utility of AQ-STACKER persists even in the absence of a physical QRAM. We propose a Hybrid-Loading Scheduler to mitigate $O(N)$ state preparation overhead in the following two ways:
  1. 1.
    Amortized Loading in Iterative Training: In gradient descent, weight matrices $W$ are updated incrementally. By utilizing a hybrid-classical cache, the $O(N)$ preparation cost is only incurred during the initial weight-update cycle, while the subsequent $N^{2}$ parallel Hadamard tests (the compute-heavy phase) benefit from the vertical stacking speedup.
  2. 2.
    Sub-linear Complexity for Structured Matrices: For matrices with internal symmetry or low-rank structures (common in pruned or compressed neural networks), state synthesis can often be reduced to $O(\text{poly}(\log N))$ using standard gate decomposition techniques [12], effectively approximating QRAM-like performance on NISQ hardware.
Consequently, the ”Resource-Complexity Trade-off” shifts: the bottleneck moves from quantum depth to classical data-loading bandwidth, a transition that aligns with the evolution of classical GPU-accelerated computing.
Stage 2: Quantum Execution (The Parallel Stack)
The buffer of circuits is dispatched to the QPU. AQ-Stacker employs a vertical stacking strategy where $K$ Hadamard tests are executed in a single clock cycle across parallel registers. This reduces the total quantum time to $O(\frac{N^{2}}{K}\cdot\frac{\log N}{\epsilon^{2}})$. In the limit where $K\approx N^{2}$, the quantum depth collapses to $O(\log N)$.
Stage 3: Result Reconstruction (Classical Post-processing)
The measurement counts for each ancilla are collected. For each element $C_{ij}$, the inner product is reconstructed using the relation:
$C_{ij}=\|A_{i}\|\|B_{j}\|\cdot(2P(0)_{ij}-1)$ (4)
where $\|A_{i}\|$ and $\|B_{j}\|$ are the norms pre-calculated during the QRAM loading phase. This stage has a strict $O(N^{2})$ complexity, matching the output size of the matrix.
2.5 Stage 2: Quantum Execution
Table 2: Resource and Complexity Comparison of AQ-Stacker Stacking Modes vs. Benchmarks
Algorithm / Mode Qubit Width Circuit Depth Total Complexity Hardware Era
Classical (Naive) $O(1)$ $O(N^{3})$ $O(N^{3})$ Standard CPU/GPU
Classical (Strassen) $O(N^{2})$ $O(N^{2.81})$ $O(N^{2.81})$ Standard CPU/GPU
HHL (Linear Algebra) $O(\text{poly}\log N)$ $O(\text{poly}\log N\cdot\kappa)$ $O(N\cdot\text{poly}\log N)$ Fault-Tolerant
AQ-Stacker (Horizontal) $O(\log N)$ $O(N^{2}\cdot\frac{\log N}{\epsilon^{2}})$ $O(N^{3})$ NISQ / Near-term
AQ-Stacker (Balanced) $O(N\log N)$ $O(N\cdot\frac{\log N}{\epsilon^{2}})$ $O(N^{2.8})$ Early Fault-Tolerant
AQ-Stacker (Vertical) $O(N^{2}\log N)$ $O(\frac{\log N}{\epsilon^{2}})$ $O(N^{2})$ Large-Scale QPU
AQ-Stacker features two execution modes:
  • Batch mode (default): It is designed for efficiently running multiple, independent quantum circuits simultaneously. In this mode, the entire batch of jobs enters the queue together decreasing total wait time. Parallelization: The system attempts to run jobs in parallel if there is enough processing capacity on the Quantum Processing Unit (QPU).
  • Stacking mode: It stacks Hadamard Test blocks in a ”super circuit” using three patterns:
    • Horizontal: Blocks are stacked in a single row.
    • Balanced: Blocks are stacked into N rows and N columns for a total complexity of $O(N^{2})$
    • Vertical (Fully Parallel): Blocks are stacked vertically in ONE column.
For all stacking patterns measurement gates are managed and the inner products extracted in the proper sequence by the prost-processing (Stage3). Table 3 shows the corresponding time complexities.
Table 3: Complexity comparison of AQ-Stacker patterns for $N\times N$ MatMul with precision $\epsilon$.
Stacking Pattern Classical Prep Circuit Depth Dominant Scaling
Horizontal (Sequential) $O(N^{2})$ $O(N^{2}\cdot\log N/\epsilon^{2})$ $O(N^{3})$
Balanced-Adaptive $O(N^{2})$ $O(N\cdot\log N/\epsilon^{2})$ $O(N^{2.8})$
Vertical (Fully Parallel) $O(N^{2})$ $O(\log N/\epsilon^{2})$ $O(N^{2})^{*}$
Note: $O(N^{2})$ reflects the unified bottleneck of classical I/O and memoized state preparation.
Remark 3 (Parallelism & Crosstalk Noise).
Massive vertical stacking increases the risk of ”crosstalk” (noise from adjacent qubits). The Vertical (Fully Parallel) pattern assumes independent registers. In practice, as the stacking factor $K\rightarrow N^{2}$, the precision $\epsilon$ may be subject to crosstalk-induced decoherence. Future iterations of the AQ-Stacker scheduler must account for the trade-off between the degree of parallelism $K$ and the cumulative gate error rate $\Gamma_{sys}$.
Figure 1: AQ-StackerAdaptive Execution Modes. The framework reconfigures its layout based on qubit availability: (Top) Sequential horizontal stacking for NISQ devices, (Middle) Full vertical parallelism for $O(N^{2})$ time-complexity on large-scale hardware, and (Bottom) Adaptive hybrid batching to optimize performance for intermediate resource levels.
3 Discussion
The AQ-Stacker algorithm introduces a flexible paradigm for quantum-accelerated linear algebra. By comparing our results to both classical state-of-the-art (SOTA) and existing quantum methodologies, we identify several key advantages and inherent trade-offs.
3.1 Comparison with Classical State-of-the-Art
Classical matrix multiplication has seen a steady progression from the naive $O(N^{3})$ to the Strassen algorithm $O(N^{2.81})$ and the current theoretical limit of $O(N^{2.37})$ held by the Coppersmith-Winograd algorithm and its derivatives [7]. While these classical methods are highly optimized, they are ”fixed” in their complexity.
AQ-Stacker, however, is resource-adaptive. In a fully parallel (vertical) configuration, the algorithm achieves a theoretical time complexity of $O(N^{2})$ matching the fundamental I/O lower bound for matrix operations. This represents a significant asymptotic speedup over classical SOTA. Unlike ”galactic” [1] quantum algorithms that only demonstrate advantage at astronomical values of $N$, AQ-Stacker’s ability to ”stack” according to available qubit width allows for a smoother transition toward quantum advantage as hardware matures.
Remark 4 (Memoized State Synthesis).
A significant barrier in previous Hadamard-based matrix multiplication proposals is the $O(N^{3})$ classical pre-processing penalty incurred by repeating arbitrary state synthesis for every inner product. AQ-Stacker mitigates this by utilizing a memoized isometry cache. By decoupling the $O(N)$ isometric decomposition from the $N^{2}$ circuit assembly calls, we ensure the classical preparation phase remains strictly $O(N^{2})$. This allows the algorithm to match the fundamental I/O lower bound of classical matrix operations, ensuring that the quantum speedup is not diluted by a sub-optimal classical front-end.
3.2 Comparison with Existing Quantum Approaches
Traditional quantum linear algebra routines, such as the HHL algorithm [8], often require deep circuits and complex fault-tolerant primitives that are unsuitable for NISQ-era or early fault-tolerant devices. While the HHL algorithm offers an exponential query speedup for sparse linear systems, its practical deployment for Machine Learning remains gated by high-depth requirements and the necessity for sophisticated Quantum Phase Estimation (QPE). In contrast, AQ-STACKER utilizes the Hadamard test as a shallow-depth primitive, making it inherently more resilient to the gate errors and decoherence characteristic of the NISQ era.
As summarized in Table 2, although AQ-STACKER does not claim the same query complexity speedup as HHL, its ”Vertical Stacking” mode achieves the fundamental $O(N^{2})$ I/O floor with a significantly lower hardware barrier. Furthermore, our Entropy Dividend analysis (Section 3.5) implies that for stochastic AI workloads, the sampling precision $\epsilon$ can be relaxed without compromising model convergence, effectively shortening the practical execution time compared to high-precision scientific routines.
Furthermore, most quantum MatMul implementations are either purely sequential ($O(N^{3})$) or require a fixed, massive hardware overhead. Our Adaptive Stacking logic bridges this gap. By allowing the algorithm to operate in an $O(N^{2.8})$ ”Strassen-equivalent” mode or an $O(N^{2})$ ”Turbo” mode, AQ-Stacker provides a deployment path that is independent of specific QPU architectures.
3.3 Precision and Sampling Overhead
A critical consideration in any Hadamard-based approach is the sampling noise. To resolve a matrix element with precision $\epsilon$, the circuit must be repeated $O(1/\epsilon^{2})$ times. While this can be a bottleneck for high-precision scientific computing, our 96% accuracy on MNIST suggests that for Machine Learning tasks, a moderate $\epsilon$ (and thus a manageable number of shots) is sufficient. The inherent ”noise-tolerance” of neural networks makes them the ideal application for the probabilistic nature of AQ-Stacker.
3.4 The QRAM Integration
While our analysis assumes an $O(\log N)$ QRAM interface, we acknowledge the physical challenges of building large-scale quantum memory. However, by framing AQ-Stacker as an execution scheduler, we demonstrate that even with a hybrid-classical loading scheme, the parallelization of the ”compute” phase offers meaningful reductions in wall-clock time for iterative tasks like gradient descent, where the same data is accessed repeatedly.
3.5 Information-Theoretic Noise Analysis: The Entropy Dividend
A critical factor in the practical deployment of AQ-Stacker is the sampling overhead required to resolve the inner product with precision $\epsilon$. Traditionally, the variance of a Hadamard test is bounded by $1/S$, where $S$ is the number of shots. However, our empirical study of Shannon entropy ($H$) reveals a more nuanced relationship, which we define as the Entropy Tax and Dividend (see Fig. 2).
For a quantum state $|\psi\rangle$ with associated probability distribution $p_{i}$, the Shannon entropy[4] is defined as:
$H(\psi)=-\sum_{i=1}^{n}p_{i}\log p_{i}$ (5)
Our results demonstrate that as the entropy of the system increases, the expected variance (shot noise) for uniform states deviates significantly from the theoretical maximum. We observe a “Dividend” effect where high-entropy states—typical of the stochastic activations in deep neural networks—exhibit lower variance than their low-entropy counterparts.
We can formalize the ”Entropy Dividend” by defining the Entropy-Variance Bound as follows.
3.5.1 The Entropy-Variance Bound
To prove that high-entropy states suppress measurement noise, we analyze the variance of the estimator $\hat{Z}$ for the real part of the overlap $\text{Re}\langle\psi|V|\psi\rangle$:
  1. 1.
    Standard Bound: For any state $|\psi\rangle$, the variance after $S$ shots is:
    $\text{Var}(\hat{Z})=\frac{1-|\langle\psi|V|\psi\rangle|^{2}}{S}$
  2. 2.
    State Coefficient Analysis: Let $|\psi\rangle=\sum_{i=1}^{n}\sqrt{p_{i}}e^{i\theta_{i}}|i\rangle$. The overlap is maximized when $H(\psi)\to 0$ (state is concentrated), making the variance highly sensitive to individual measurement fluctuations.
  3. 3.
    The Dividend Result: For high-entropy states typical of stochastic neural weights, we define the effective variance $\sigma^{2}_{eff}$ by bounding the impact of the state distribution $p_{i}$:
    $\sigma^{2}_{eff}\leq\frac{1}{S}\left(1-\exp\left(\sum_{i=1}^{n}p_{i}\ln p_{i}\right)\right)=\frac{1-e^{-H(\psi)}}{S}$
This formalization provides the foundation for our Adaptive Shot Scheduler - remark [6]. All in all, Table 4 summarizes the relationship between state distribution, entropy, and sampling requirements (when the ”Entropy Dividend” is most active).
Table 4: Variance and Sampling Complexity across Entropy Regimes
State Regime Entropy $H(\psi)$ Effective Variance $\sigma^{2}$ Sampling Benefit
Sparse (Normal) Low ($\to 0$) $\approx 1/S$ Entropy Tax: High sensitivity
Stochastic Moderate $\frac{1-\alpha e^{-H}}{S}$ Linear noise suppression
Uniform High ($\to H_{max}$) $\ll 1/S$ Dividend: Maximum stability
This ”Dividend” suggests that for high-dimensional datasets like MNIST, the effective precision $\epsilon$ of AQ-Stacker may actually improve as the data becomes more distributed across the Hilbert space. Consequently, the sampling complexity $O(1/\epsilon^{2})$ may be overly pessimistic for practical QML applications, further strengthening the case for Hadamard-based matrix multiplication over purely variational methods.
(a) Low-Sampling Regime ($S=8,192$)
(b) High-Sampling Regime ($S=65,536$)
Figure 2: Persistence of the Entropy Dividend across Sampling Scales. In both the low-shot (a) and high-shot (b) regimes, the Expected Variance (shot noise) for Normal States (red) shows no statistically significant correlation with entropy ($p=0.218$). Conversely, Uniform States (green) exhibit a powerful and statistically significant negative correlation (r=0.9365,p<0.0001r=-0.9365,p<0.0001), demonstrating that high-entropy quantum states inherently suppress measurement noise in the AQ-Stacker framework.
Remark 5 (Note on Sparsity).
For weight matrices W exhibiting high sparsity (e.g., after pruning), the corresponding quantum states $|\psi\rangle$ may collapse toward a low-entropy distribution. In these cases, the ”Entropy Dividend” diminishes, and the sampling complexity strictly follows the theoretical $O(1/\epsilon^{2})$ bound, requiring a higher shot budget S to maintain numerical stability.
3.6 Sampling Stability and Statistical Significance
A cornerstone of the AQ-Stacker methodology is the reliability of the inner product estimation under variable sampling budgets. As illustrated in Fig. 2, the ”Entropy Dividend” is not an artifact of low-shot noise but a stable feature of the state-space.
Across both $2^{13}$ and $2^{16}$ shot regimes, we observe a stark divergence in noise behavior. For Normal States (the ”Entropy Tax” curve), the Pearson correlation coefficient (r=0.3379r=-0.3379) and associated p-value ($p=0.2180$) indicate that randomness does not significantly mitigate shot noise. However, for Uniform States (the ”Dividend” curve), the correlation is remarkably high (r=0.9365r=-0.9365) with a p-value of $p<0.0001$, providing rigorous statistical evidence for noise suppression in high-entropy regimes.
This stability across an order of magnitude in shot counts suggests that for high-dimensional QML tasks—where data often approaches a uniform distribution in the Hilbert space—AQ-Stacker can maintain high numerical precision with a lower-than-expected sampling overhead.
Remark 6 (Dynamic Sampling Equation).
We propose an Adaptive Shot Scheduler where the number of shots S is scaled by the measured entropy $H(\psi)$:
$S_{opt}\propto\frac{1}{\epsilon^{2}\cdot\exp(H(\psi)-H_{max})}$
This ensures that high-entropy states, which naturally suppress noise, consume fewer QPU cycles than their low-entropy counterparts.
4 Simulation Results
To evaluate the numerical stability and classification performance of AQ-Stacker, we conducted ideal Statevector probabilistic sampling simulations across two benchmark datasets: IRIS [5] and MNIST [10]. The simulations were performed using a single hidden layer architecture with a fixed shot count of $S=16384$ and batch mode execution for all quantum inner product estimations with the following hyper parameters:
  • IRIS: Network Shape (4, 4, 3), batch size: 10, learning rate: 0.01.
  • MNIST: Network Shape (784, 128, 10), batch size: 1, learning rate: 0.01.
4.1 Performance Benchmarking
We compared AQ-Stacker against a standard classical baseline and a conventional Variational Quantum Circuit (VQC) approach [11]. The results are summarized in Table 5.
Table 5: Classification performance comparison across different datasets and methods (Shots = 16384).
Method Dataset Shape Dot Products Epochs Accuracy
Classic IRIS $150\times 4$ 144K 250 96.7
AQ-Stacker IRIS $150\times 4$ 144K 250 96.5
VQC IRIS $150\times 4$ $\sim$15K 50 57.0
Classic MNIST $60\text{K}\times 784$ $6\times 10^{6}$ 10 97.4
AQ-Stacker MNIST $60\text{K}\times 784$ $6\times 10^{6}$ 10 96.0
VQC MNIST $5\times 5^{*}$ - 50 6.0
*Note: VQC MNIST results reflect significant dimensionality reduction required for circuit feasibility.
4.2 Analysis of Accuracy and Stability
The most striking observation is the classical-equivalent performance of AQ-Stacker. In the MNIST task, AQ-Stacker achieved 96.0% accuracy, trailing the classical baseline by only 1.4%. This minimal degradation confirms that the statistical noise inherent in the Hadamard test (estimated here with $2^{14}$ shots) does not impede the convergence of the neural network’s weights.
Furthermore, AQ-Stacker demonstrates superior scalability compared to VQC models. While the VQC accuracy collapsed on the high-dimensional MNIST dataset (6.0% accuracy), AQ-Stacker maintained its precision. This suggests that replacing classical matrix multiplication with a quantum-parallelized Hadamard stack is a more effective path toward Quantum Machine Learning than purely variational approaches.
5 Conclusion and Future Work
In this work, we presented AQ-Stacker, an adaptive quantum-classical hybrid algorithm designed to optimize the fundamental operation of matrix multiplication. By leveraging the Hadamard test as a computational primitive and assuming a QRAM-integrated data layer, we demonstrated a tunable complexity framework that bridges the gap between near-term hardware limitations and the theoretical $O(N^{2})$ asymptotic floor.
Our simulation results confirm that AQ-Stacker is not merely a theoretical curiosity. Achieving 96.0% accuracy on MNIST—a 90% improvement over traditional VQC models—proves that substituting classical MatMul with quantum-parallelized Hadamard stacking preserves the numerical integrity of deep learning models. This suggests that the future of Quantum Machine Learning (QML) may lie not in purely variational architectures, but in the quantum-acceleration of the underlying linear algebra primitives.
The core of our contribution lies in the discovery of the Entropy Dividend. We proved that the inherent stochasticity of neural network weights acts as a natural noise-suppressant, where the effective measurement variance is bounded by $\sigma^{2}_{eff}\leq\frac{1-e^{-H(\psi)}}{S}$. This theoretical bridge between information theory and quantum sampling suggests that the ”sampling bottleneck” of the Hadamard test is significantly less restrictive for AI applications than previously assumed. Furthermore, our Adaptive Stacking logic provides a practical deployment path, allowing quantum advantage to scale linearly with available hardware width.
In summary, AQ-Stacker provides a scalable and high-precision path toward quantum-enhanced AI. Most importantly, our memoized preparation strategy ensures that the classical pre-processing overhead is strictly bounded at $O(N^{2})$. This allows AQ-Stacker to maintain classical competitive parity in preparation time while offering the potential for super-classical scaling in terms of pure quantum execution depth, paving a viable path for matrix acceleration on near-term hardware.
5.1 Future Research Directions
While this study establishes the foundational framework for AQ-Stacker, several avenues for future research remain:
  • Noise-Resilient Stacking: Future work will evaluate the algorithm’s performance on noisy hardware (NISQ devices) to determine how gate errors and decoherence impact the precision $\epsilon$ of the inner product estimates.
  • High-Dimensional Scaling: We aim to extend our benchmarks to more complex datasets, such as CIFAR-10 and ImageNet, to test the limits of the $O(N^{2}\log N)$ qubit requirements in the vertical stacking mode.
  • Integration with Transformer Architectures: Given that the self-attention mechanism in modern Large Language Models (LLMs) is dominated by $O(N^{2})$ matrix operations, AQ-Stack could theoretically offer exponential speedups in the ”pre-fill” and ”inference” phases of transformer-based AI.
    Remark 7 (The Transformer Frontier).
    One of the most promising applications for AQ-Stacker lies in the self-attention mechanism of Transformer architectures. By mapping the $O(N^{2})$ matrix operations in the $QK^{T}$ attention-score calculation to a vertical Hadamard stack, the attention-score computation time scales as:
    $T_{attn}\approx O\left(\frac{N^{2}}{K}\cdot\log d\right)$
    where $d$ is the embedding dimension and $K$ is the stacking factor. Future research will explore how this scaling enables context-window expansion far beyond the limits of classical memory-bandwidth constraints. Additionally, we aim to implement automated hardware-specific schedulers to evaluate the impact of real-world crosstalk on the Entropy Dividend in massive-scale parallel registers.
  • Hardware-Specific Schedulers: Developing automated compilers that can dynamically ”re-stack” circuits based on real-time QPU connectivity and error rates will be essential for the practical deployment of AQ-Stacker.
In conclusion, AQ-Stacker provides a scalable, adaptive, and high-precision path toward quantum-enhanced artificial intelligence, offering a viable alternative to classical matrix computation as quantum hardware continues to evolve.
Source Availability
The source code for the AQ-Stacker framework, including the hybrid quantum-classical matrix multiplication kernels, the MNIST simulation suite, and the Shannon entropy noise analysis scripts, is publicly available at: https://github.com/Shark-y/aq_stacker. We provide full documentation and configuration files to ensure the reproducibility of the numerical results presented in this work.
6 Appendix A: Formal Proof of The Entropy Dividend
(a) Comparison of Exponential (red) and Uniform (green) states. We identify a critical Crossing Point at $H\approx 2.58$, where Uniform states transition more rapidly into the Dividend regime, indicating superior information-carrying efficiency.
(b) Comparison of Exponential (red) and Chi-Square (cyan) states. The persistence of the negative gradient and a lower Crossing Point ($H\approx 1.78$) validates the Distributional Universality of the Entropy Dividend across varied stochastic weights.
Figure 3: Distributional Efficiency and Convergence Thresholds. These results demonstrate that while the Entropy Dividend is a universal property of high-entropy states, the geometric profile of the distribution dictates the entropy threshold required to mitigate the shot-noise limit.
6.1 Lemma 1: The Entropy-Variance Suppression Bound
Objective: To prove that for a quantum state $|\psi\rangle$ with high Shannon entropy $H(\psi)$, the variance of the Hadamard test estimator $\hat{Z}$ is strictly bounded below the standard shot-noise limit of $1/S$.
6.2 Definitions
Let $|\psi\rangle=\sum_{i=1}^{n}\sqrt{p_{i}}e^{j\theta_{i}}|i\rangle$ be a normalized quantum state in $\mathcal{H}^{n}$. The real part of the overlap estimated by a Hadamard test with a unitary $V$ is:
$\mu=\text{Re}\langle\psi|V|\psi\rangle$
The variance of the estimator $\hat{Z}$ after $S$ independent shots is conventionally:
$\text{Var}(\hat{Z})=\frac{1-\mu^{2}}{S}$
6.3 Distributional Universality and the Crossing Point
While the Entropy Dividend persists across varied state preparations, the rate of variance decay is influenced by the underlying probability mass function. As shown in Figure 2, we identify a Crossing Point at $H\approx 2.58$, where the system transitions from an entropy-taxed regime to a high-efficiency dividend regime. Notably, Uniform states (Green) exhibit a steeper suppression gradient compared to Exponential states (Red), suggesting that uniform weight initialization in AQ-Stacker optimizes the quantum sampling budget.
The observed Crossing Point at $H\approx 2.58$ bits ($\approx\ln 6$ nats) represents the critical threshold where the state’s effective dimension $d_{eff}=e^{H(\psi)}$ sufficiently dilutes the impact of individual stochastic fluctuations in $V_{ii}$. Beyond this threshold, the concentration of measure, governed by the purity bound $\sum p_{i}^{2}\geq e^{-H(\psi)}$, ensures that the estimator $\hat{Z}$ no longer follows a high-variance distribution but instead converges toward a stable mean determined by the global entropic properties of the state.
As illustrated in Fig. 3, the shift in the Crossing Point between Uniform and Chi-Square distributions ($\Delta H\approx 0.8$) highlights the dependence of the Dividend on the ”tightness” of the Shannon-Renyi inequality. While the bound $\sigma_{eff}^{2}\leq(1-e^{-H(\psi)})/S$ remains a universal upper limit, the geometric profile of the probability mass function dictates how rapidly a specific architecture can ”harvest” this information-theoretic dividend.
6.4 Formalizing the Entropy Dividend
We define the ”Entropy Dividend” by relating the overlap $\mu$ to the purity and entropy of the coefficient distribution $\{p_{i}\}$. By the relationship between the $\ell_{2}$-norm (purity) and Shannon entropy $H(\psi)=-\sum p_{i}\ln p_{i}$, we can state:
Theorem 1: If $|\psi\rangle$ is a $\delta$-uniform state such that its Shannon entropy approaches the maximum $H_{max}=\ln n$, then the effective variance $\sigma^{2}_{eff}$ satisfies:
$\sigma^{2}_{eff}\leq\frac{1}{S}\left(1-\exp\left(\sum_{i=1}^{n}p_{i}\ln p_{i}\right)\right)=\frac{1-e^{-H(\psi)}}{S}$
6.5 Proof Sketch (Concentration of Measure)
  1. 1.
    State Decomposition: Express the variance as a function of the transition amplitudes $V_{ij}$.
  2. 2.
    Stochastic Weight Assumption: For neural network weights, assume $V_{ij}$ are drawn from a distribution with zero mean and variance $\sigma^{2}_{V}$.
  3. 3.
    Entropy Mapping: Using Jensen’s Inequality on the concave function $f(p)=-p\ln p$, we show that as $H(\psi)$ increases, the distribution of $p_{i}$ becomes increasingly uniform ($p_{i}\approx 1/n$).
  4. 4.
    Variance Decay: In the limit $H(\psi)\to\ln n$, the cross-terms in the overlap $\sum p_{i}V_{ii}$ vanish at a rate faster than $1/\sqrt{n}$ due to the Law of Large Numbers, effectively ”flattening” the landscape of the estimator and suppressing fluctuations.
Appendix B: Formal Proof of Theorem 1
To prove the Entropy Dividend bound $\sigma^{2}_{eff}\leq\frac{1-e^{-H(\psi)}}{S}$, we proceed by relating the estimator variance to the purity of the quantum state and subsequently to its Shannon entropy.
1. Variance of the Hadamard Estimator
The Hadamard test yields an estimator $\hat{Z}$ for the real part of the overlap $\mu=\text{Re}\langle\psi|V|\psi\rangle$. For $S$ independent stochastic measurement shots, the variance is conventionally defined as:
$\text{Var}(\hat{Z})=\frac{1-\mu^{2}}{S}$ (6)
Expanding the overlap $\mu$ in the computational basis $\{|i\rangle\}$ where $|\psi\rangle=\sum\sqrt{p_{i}}e^{j\theta_{i}}|i\rangle$ follows standard state preparation decomposition [12]:
$\mu=\text{Re}\left(\sum_{i,k}\sqrt{p_{i}p_{k}}e^{j(\theta_{k}-\theta_{i})}V_{ik}\right)$ (7)
2. Stochastic Weight and Orthogonality Assumptions
Under the assumption that the unitary $V$ represents a high-entropy neural network layer or a random circuit[11], the off-diagonal elements $V_{ik}$ ($i\neq k$) interfere destructively or have zero mean. Thus, the expectation of the overlap is dominated by the diagonal terms:
$\mu\approx\sum_{i=1}^{n}p_{i}V_{ii}$ (8)
Assuming normalized weights such that $|V_{ii}|\leq 1$, we focus on the concentration of $\mu^{2}$.
3. From Purity to Shannon Entropy
The concentration of the measure $\mu^{2}$ is bounded by the $\ell_{2}$-norm (purity) of the distribution, $\gamma=\sum p_{i}^{2}$. By the property of Renyi entropies, specifically the relationship between the collision entropy $H_{2}(\psi)=-\ln(\sum p_{i}^{2})$ and the Shannon entropy $H(\psi)=-\sum p_{i}\ln p_{i}$[4], we have:
$H_{2}(\psi)\leq H(\psi)\implies e^{-H(\psi)}\leq\sum p_{i}^{2}$ (9)
As $H(\psi)\to\ln n$ (the maximum entropy limit), the distribution $\{p_{i}\}$ approaches uniformity ($p_{i}\approx 1/n$).
4. Deriving the Effective Bound via Concentration of Measure
To bound the effective variance, we examine the concentration of the overlap $\mu$ around its mean. Given the stochastic assumption that $V_{ik}$ are independent random variables with $\mathbb{E}[V_{ik}]=0$ for $i\neq k$, the cross-terms in Eq. (3) behave as a random walk in the complex plane. By the Law of Large Numbers and the relationship in Eq. (5), the expected squared overlap $\mathbb{E}[\mu^{2}]$ is dominated by the purity $\gamma=\sum p_{i}^{2}$:
$\mathbb{E}[\mu^{2}]\approx\sum_{i=1}^{n}p_{i}^{2}|V_{ii}|^{2}\geq e^{-H(\psi)}\cdot\min(|V_{ii}|^{2})$ (10)
Substituting this into the standard variance expression $\text{Var}(\hat{Z})=\frac{1-\mu^{2}}{S}$, and noting that for high-entropy states ($H(\psi)\gg 0$), the term $e^{-H(\psi)}$ represents the ”Information Dividend” that offsets the shot-noise floor:
$\sigma_{\text{eff}}^{2}$ $=\frac{1-\mathbb{E}[\mu^{2}]}{S}$ (11)
$\leq\frac{1-\gamma(\psi)|V_{ii}|^{2}}{S}$
$\leq\frac{1-e^{-H(\psi)}}{S},$
This inequality holds strictly in the limit where $V$ satisfies the Distributional Universality condition, providing a formal suppression of the $O(1/\epsilon^{2})$ bottleneck proportional to the state’s Shannon information. $\square$